OpenStudy (anonymous):

A twelve-sided die has sides numbered 1 through 12. The die is rolled once. Find each probability. with and without replacement

OpenStudy (anonymous):

17. P(4 or 5) 18. P(even or a multiple of 3) 19. P(odd or a multiple 4)

OpenStudy (anonymous):

one question nancy: why with or without replacement if die is only rolled once?

OpenStudy (anonymous):

if one roll then 17. P = 1/6 18 P = (1/2 + 1/3) = 5/6 19. P = (1/2 + 1/4) = 3/4